You are listening to the siren of an emergency vehicle that is stuck in a traffic jam. Sound from the siren of the vehicle has a frequency of 703.4 Hz and the speed of sound waves on this day is 335.2 m/s. (a) What is the distance from one condensation to the next? Give your answer in meters. m (b) What is the distance from one condensation to the nearest rarefaction?
Answer #1

(a) What is the distance from one condensation to the next?
v = fλ and λ = v/f
v = 335.2 m/s f = 703.4 Hz

λ = (335.2 m/s) / (703.4/s)
= 0.4765 m

b)   The second question is asking you for half that distance because you're only going from the high peak to the nearest low point. The low point is right between the two high peaks. Therefore, you only have to divide your first answer by two.

0.4765m / 2 = .2382 m